The amount of paint needed to cover the walls of a room varies jointly as the perimeter of the room and the heightof the walls. If a room with a perimeter of 40 feet and a 10 foot walls takes 4 quarts of paint, find the amount of paint needed to cover 80 feet and 8 foot walls.
Answer by rfer(16322)

40*10=400
100 sq ft per qt
80*8=640/100=6.4 qts